# October 17, 2025

### v2

@@ IV Mathematica notebook
To optimize this process and reduce the burden on the high-energy physics phenomenology community, we have developed the Mathematica tool BounDS that automates the calculation. The user simply has to specify the values of $n_{D}$, $n_{c}$, and $n_{n}$, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries. BounDS carries out the necessary steps to:
@@
V Unitarity Bounds for particular cases
We now proceed to study unitarity bounds in specific scenarios by specifying the values of $n_{D}$, $n_{c}$, and $n_{n}$, as well as any additional symmetries. The cases studied explicitly in this work are summarized in Table 2.
@@ V Unitarity Bounds for particular cases
2 0 0 $\mathbb{Z}_{2}$ Ginzburg:2005dt
2 0 0 Ginzburg:2005dt
1 2 0 Muhlleitner:2020wwk generalized
@@
2 2 0 $\mathbb{Z}_{2}\otimes\mathbb{Z}^{\prime}_{2}$ Boto:2025mmn
@@
3 0 0 $\mathbb{Z}_{3}$ Bento:2017eti
Table 2: Models presented as examples in this article.
@@ V.1.1 Scattering Matrices
V.1.2 Unitarity Bounds
Applying the partial wave unitarity condition of Eq. (24) to the eigenvalues of these scattering matrices, we find
@@ V.1.2 Unitarity Bounds
In order to facilitate the comparison of our results with those of Ref. Bento:2017eti , it is useful to provide a short dictionary of notations. Table 3 lists the correspondence between the coupling employed in this work and the one used in Ref. Bento:2017eti .
Term Our Notation Notation in Bento:2017eti
@@
Table 3: Comparison of coupling notation.
This recovers the classic results of Lee:1977eg .
@@ V.2.1 Scattering Matrices
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&2\sqrt{2}\gamma_{11,11}\\ 2\gamma_{11,11}&12\beta_{11,11}\\ \end{bmatrix}$. (51)
V.2.2 Unitarity Bounds
We apply the partial wave unitarity condition of Eq. (24) to each of the eigenvalues of the zero partial wave amplitude matrix and find
@@ V.3 The $\mathbb{Z}_{2}$-Symmetric 2HDM
where, in Eq. (58) we have employed the standard notation of Ref. Ginzburg:2005dt and, without loss of generality, choose $\lambda_{5}$ to be real. The mapping between the two notations is given in Table 4.
Term Our Notation Notation in Ginzburg:2005dt
@@ V.3.2 Eigenvalues
By imposing the unitarity condition of Eq. (24) to each of the eigenvalues of the zero partial wave amplitude matrix we derive
@@
We find that our results are in agreement with those derived in Ginzburg:2005dt .
@@ V.4 The general 2HDM
where, in Eq. (70) we have employed, once again, the standard notation of Ref. Ginzburg:2005dt . The mapping between the two notations is given in Table 5.
Term Our Notation Notation in Ginzburg:2005dt
@@ V.4.1 Scattering Matrices
$M_{|2,1,1\rangle}=\begin{bmatrix}\lambda_{1}&\sqrt{2}\,\lambda_{6}^{*}&\lambda_{5}^{*}\\ \sqrt{2}\,\lambda_{6}&\lambda_{3}+\lambda_{4}&\sqrt{2}\,\lambda_{7}^{*}\\ \lambda_{5}&\sqrt{2}\,\lambda_{7}&\lambda_{2}\end{bmatrix}$, (71)
$M_{|1,0,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{6}&\lambda_{6}^{*}&\lambda_{4}\\ \lambda_{6}^{*}&\lambda_{3}&\lambda_{5}^{*}&\lambda_{7}^{*}\\ \lambda_{6}&\lambda_{5}&\lambda_{3}&\lambda_{7}\\ \lambda_{4}&\lambda_{7}&\lambda_{7}^{*}&\lambda_{2}\end{bmatrix}$, (72)
@@
Our scattering matrices agree with those in Ginzburg:2005dt . If we set $\lambda_{6}=\lambda_{7}=0$ (along with their complex conjugates), we re-obtain, as expected, the results presented in Section V.3.1.
V.4.2 Unitarity Bounds
The eigenvalues of the scattering matrices in this case are, in general, too complex to write down in closed form. However, some can be computed analytically, and unitarity bounds are imposed accordingly,
@@ V.5.1 Scattering Matrices
$M_{|0,0,0\rangle}=\begin{bmatrix}6\,\lambda_{11,11}&2\sqrt{2}\,\gamma_{11,11}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{11,22}\\ 2\,\gamma_{11,11}&12\,\beta_{11,11}&12\sqrt{2}\,\beta_{11,12}&12\,\beta_{11,22}\\ 2\sqrt{2}\,\gamma_{11,12}&12\sqrt{2}\,\beta_{11,12}&24\,\beta_{11,22}&12\sqrt{2}\,\beta_{12,22}\\ 2\,\gamma_{11,22}&12\,\beta_{11,22}&12\sqrt{2}\,\beta_{12,22}&12\,\beta_{22,22}\end{bmatrix}$. (79)
V.5.2 Unitarity Bounds
Again, certain eigenvalues of the scattering matrices are too complicated to be expressed in closed form. However, others can be written analytically, and, for those, the corresponding unitarity bounds are then applied, as:
@@ V.5.2 Unitarity Bounds
In order to compare our results with previous work, we consider the model studied in Ref. Muhlleitner:2020wwk , where the SM is extended by a neutral complex scalar singlet. In their notation, the quartic part of the scalar potential reads
@@
where  [^5] Equivalently, one could choose $S=\frac{\chi_{2}+i\chi_{1}}{\sqrt{2}}$. $S=\frac{\chi_{1}+i\chi_{2}}{\sqrt{2}}$. Notice that the case with two real singlets in Eq. (76) is more general than the case with one complex singlet in Eq. (83). But, since the complex scalar singlet can be decomposed into two real scalar singlet fields, $\chi_{1}$ and $\chi_{2}$, we find that our results reproduce those of Muhlleitner:2020wwk , when we set $\gamma_{11,12}=\beta_{11,12}=\beta_{12,22}=0$. By matching the terms in Eq. (83) to those in Eq. (76), we derive the mapping between the two notations that is provided in Table 6.
Term Our Notation Notation in Muhlleitner:2020wwk
@@ V.6.1 Scattering Matrices
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda^{*}_{12,12}&0\\ 2\lambda_{12,12}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (87)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda^{*}_{12,12}\\ 0&0&2\lambda_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (88)
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&4\lambda_{11,22}+2\lambda_{12,21}&2\sqrt{2}\,\gamma_{11,11}&0&0\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda_{22,22}&2\sqrt{2}\,\gamma_{22,11}&0&0\\ 2\gamma_{11,11}&2\gamma_{22,11}&12\beta_{11,11}&0&0\\ 0&0&0&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,12}\\ 0&0&0&6\lambda^{*}_{12,12}&2\lambda_{11,22}+4\lambda_{12,21}\end{bmatrix}$, (89)
@@ V.6.2 Unitarity Bounds
$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq 8\pi.$ (96)
@@ V.7.1 Scattering Matrices
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}2\gamma_{11,11}&2\gamma_{11,12}&0&0\\ 2\gamma_{11,12}&2\gamma_{11,22}&0&0\\ 0&0&2\gamma_{22,11}&2\gamma_{22,12}\\ 0&0&2\gamma_{22,12}&2\gamma_{22,22}\end{bmatrix}$, (100)
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda^{*}_{12,12}&0\\ 2\lambda_{12,12}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (101)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda^{*}_{12,12}\\ 0&0&2\lambda_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (102)
@@
$A=\begin{bmatrix}24\,\beta_{11,22}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{22,12}&12\sqrt{2}\,\beta_{11,12}&12\sqrt{2}\,\beta_{12,22}\\ 2\sqrt{2}\,\gamma_{11,12}&6\,\lambda_{11,11}&4\,\lambda_{11,22}+2\,\lambda_{12,21}&2\sqrt{2}\,\gamma_{11,11}&2\sqrt{2}\,\gamma_{11,22}\\ 2\sqrt{2}\,\gamma_{22,12}&4\,\lambda_{11,22}+2\,\lambda_{12,21}&6\,\lambda_{22,22}&2\sqrt{2}\,\gamma_{22,11}&2\sqrt{2}\,\gamma_{22,22}\\ 12\sqrt{2}\,\beta_{11,12}&2\,\gamma_{11,11}&2\,\gamma_{22,11}&12\,\beta_{11,11}&12\,\beta_{11,22}\\ 12\sqrt{2}\,\beta_{12,22}&2\,\gamma_{11,22}&2\,\gamma_{22,22}&12\,\beta_{11,22}&12\,\beta_{22,22}\end{bmatrix}$, (105)
$B=\begin{bmatrix}2\,\lambda_{11,22}+4\,\lambda_{12,21}&6\,\lambda_{12,12}\\ 6\,\lambda_{12,12}^{*}&2\,\lambda_{11,22}+4\,\lambda_{12,21}\end{bmatrix}$. (106)
V.7.2 Unitarity Bounds
@@
$\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{\gamma_{11,11}^{2}+4\gamma_{11,12}^{2}-2\gamma_{11,11}\gamma_{11,22}+\gamma_{11,22}^{2}}\right|$ $\leq$ $8\pi,$ (112)
$\left|\gamma_{22,11}+\gamma_{22,22}\pm\sqrt{\gamma_{22,11}^{2}+4\gamma_{22,12}^{2}-2\gamma_{22,11}\gamma_{22,22}+\gamma_{22,22}^{2}}\right|$ $\leq$ $8\pi.$ (113)
Term Our Notation Notation of Boto:2025mmn
@@
Table 7: Comparison of coupling notation.
We verified our results against those reported in Boto:2025mmn and found them to be in agreement. Table 7 summarizes the correspondence between the couplings used in our work and those of Boto:2025mmn .
@@ V.8.1 Scattering Matrices
$M_{|2,\frac{3}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}^{*}\\ \delta_{12,11}&\delta_{22,11}\end{bmatrix}$, (116)
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}&0&-2i\kappa^{*}_{12,11}\\ \delta^{*}_{12,11}&\delta_{22,11}&2i\kappa^{*}_{12,11}&0\\ 0&-2i\kappa_{12,11}&2\gamma_{11,11}&2\gamma^{*}_{12,11}\\ 2i\kappa_{12,11}&0&2\gamma_{12,11}&2\gamma_{22,11}\end{bmatrix}$, (117)
$M_{|1,1,0\rangle}=\begin{bmatrix}2\lambda_{11,22}-2\lambda_{12,21}&2i\sqrt{2}\kappa^{*}_{12,11}\\ -2i\sqrt{2}\kappa^{*}_{12,11}&2\zeta_{11,11}\end{bmatrix}$, (118)
$M_{|1,1,2\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{11,12}^{*}&2\lambda_{12,12}^{*}\\ 2\sqrt{2}\lambda_{11,12}&2\lambda_{11,22}+2\lambda_{12,21}&2\sqrt{2}\lambda_{12,22}^{*}\\ 2\lambda_{12,12}&2\sqrt{2}\lambda_{12,22}&2\lambda_{22,22}\end{bmatrix}$, (119)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{11,12}&2\lambda_{11,12}^{*}&2\lambda_{12,21}\\ 2\lambda_{11,12}^{*}&2\lambda_{11,22}&2\lambda_{12,12}^{*}&2\lambda_{12,22}^{*}\\ 2\lambda_{11,12}&2\lambda_{12,12}&2\lambda_{11,22}&2\lambda_{12,22}\\ 2\lambda_{12,21}&2\lambda_{12,22}&2\lambda_{12,22}^{*}&2\lambda_{22,22}\end{bmatrix}$, (120)
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&6\lambda^{*}_{11,12}&6\lambda_{11,12}&4\lambda_{11,22}+2\lambda_{12,21}&\sqrt{2}\delta_{11,11}&2\sqrt{2}\gamma_{11,11}\\ 6\lambda_{11,12}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda_{12,12}&6\lambda_{12,22}&\sqrt{2}\delta_{12,11}&2\sqrt{2}\gamma_{12,11}\\ 6\lambda^{*}_{11,12}&6\lambda^{*}_{12,12}&2\lambda_{11,22}+4\lambda_{12,21}&6\lambda^{*}_{12,22}&\sqrt{2}\delta^{*}_{12,11}&2\sqrt{2}\gamma^{*}_{12,11}\\ 4\lambda_{11,22}+2\lambda_{12,21}&6\lambda^{*}_{12,22}&6\lambda_{12,22}&6\lambda_{22,22}&\sqrt{2}\delta_{22,11}&2\sqrt{2}\gamma_{22,11}\\ \sqrt{2}\delta_{11,11}&\sqrt{2}\delta^{*}_{12,11}&\sqrt{2}\delta_{12,11}&\sqrt{2}\delta_{22,11}&4\alpha_{11,11}&\sqrt{2}\zeta_{11,11}\\ 2\gamma_{11,11}&2\gamma^{*}_{12,11}&2\gamma_{12,11}&2\gamma_{22,11}&\sqrt{2}\zeta_{11,11}&12\beta_{11,11}\end{bmatrix}$. (121)
V.8.2 Unitarity Bounds
Since certain eigenvalues are too complex to evaluate analytically, unitarity constraints are shown only for the remaining eigenvalues of the zero partial-wave amplitude matrix. These read
$\left|\zeta_{11,11}+\lambda_{11,22}-\lambda_{12,21}\pm\sqrt{\left(-\zeta_{11,11}-\lambda_{11,22}+\lambda_{12,21}\right)^{2}-4\left(\zeta_{11,11}\lambda_{11,22}-\zeta_{11,11}\lambda_{12,21}-2|\kappa_{12,11}|^{2}\right)}\right|$ $\leq 8\pi,$ (122)

### v3

@@ IV Mathematica notebook
To optimize this process and reduce the burden on the high-energy physics phenomenology community, we have developed the Mathematica tool BounDS that automates the calculation. The user simply has to specify the values of $n_{D}$, $n_{c}$, and $n_{n}$, and, if needed, the transformation properties of the fields under additional flavour and/or generalized CP symmetries which may either be discrete or continuous. BounDS carries out the necessary steps to:
@@
V Perturbative Unitarity Bounds for particular cases
We now proceed to study perturbative unitarity bounds in specific scenarios by specifying the values of $n_{D}$, $n_{c}$, and $n_{n}$, as well as any additional symmetries. The cases studied explicitly in this work are summarized in Table 2.
Table 2: List of models presented as examples in this article.
@@ V Perturbative Unitarity Bounds for particular cases
2 0 0 $\mathbb{Z}_{2}$ Ginzburg and Ivanov (2005)
2 0 0 Ginzburg and Ivanov (2005)
1 2 0 Mühlleitner et al. (2022) generalized
@@
2 2 0 $\mathbb{Z}_{2}\otimes\mathbb{Z}^{\prime}_{2}$ Boto et al. (2025)
@@
3 0 0 $\mathbb{Z}_{3}$ Bento et al. (2017)
@@ V.1.1 Scattering Matrices
V.1.2 Perturbative Unitarity Bounds
Applying the partial-wave unitarity condition of Eq. (26) to the eigenvalues of these scattering matrices, immediately leads to
@@ V.1.2 Perturbative Unitarity Bounds
In order to facilitate the comparison of our results with those of Ref. Bento et al. (2017), it is useful to provide a short dictionary of notations. Table 3 lists the correspondence between the coupling employed in this work and the one used in Ref. Bento et al. (2017).
Table 3: Comparison of coupling notation.
Term Our Notation Notation in Bento et al. (2017)
@@
This recovers the classic results of Lee et al. (1977b).
@@ V.2.1 Scattering Matrices
$M_{|0,0,0\rangle}=\begin{bmatrix}6\lambda_{11,11}&2\gamma_{11,11}\\ 2\gamma_{11,11}&12\beta_{11,11}\\ \end{bmatrix}$. (53)
V.2.2 Perturbative Unitarity Bounds
We apply the partial-wave unitarity condition of Eq. (26) to each of the eigenvalues of the zero partial-wave amplitude matrix and find
@@ V.3 The $\mathbb{Z}_{2}$-Symmetric 2HDM
where, in Eq. (60) we have employed the standard notation of Ref. Ginzburg and Ivanov (2005) and, without loss of generality, choose $\lambda_{5}$ to be real. The mapping between the two notations is given in Table 4.
Table 4: Comparison of coupling notation.
Term Our Notation Notation in Ginzburg and Ivanov (2005)
@@ V.3.2 Eigenvalues
By imposing the unitarity condition of Eq. (26) to each of the eigenvalues of the zero partial-wave amplitude matrix we derive
@@
We find that our results are in agreement with those derived in Ginzburg and Ivanov (2005).
@@ V.4 The general 2HDM
where, in Eq. (72) we have employed, once again, the standard notation of Ref. Ginzburg and Ivanov (2005). The mapping between the two notations is given in Table 5.
Table 5: Comparison of coupling notation.
Term Our Notation Notation in Ginzburg and Ivanov (2005)
@@ V.4.1 Scattering Matrices
$M_{|2,1,1\rangle}=\begin{bmatrix}\lambda_{1}&\sqrt{2}\,\lambda_{6}&\lambda_{5}\\ \sqrt{2}\,\lambda_{6}^{*}&\lambda_{3}+\lambda_{4}&\sqrt{2}\,\lambda_{7}\\ \lambda_{5}^{*}&\sqrt{2}\,\lambda_{7}^{*}&\lambda_{2}\end{bmatrix}$, (73)
$M_{|1,0,1\rangle}=\begin{bmatrix}\lambda_{1}&\lambda_{6}^{*}&\lambda_{6}&\lambda_{4}\\ \lambda_{6}&\lambda_{3}&\lambda_{5}&\lambda_{7}\\ \lambda_{6}^{*}&\lambda_{5}^{*}&\lambda_{3}&\lambda_{7}^{*}\\ \lambda_{4}&\lambda_{7}^{*}&\lambda_{7}&\lambda_{2}\end{bmatrix}$, (74)
@@
Our scattering matrices agree with those in Ginzburg and Ivanov (2005). If we set $\lambda_{6}=\lambda_{7}=0$ (along with their complex conjugates), we re-obtain, as expected, the results presented in Section V.3.1.
V.4.2 Perturbative Unitarity Bounds
The eigenvalues of the scattering matrices in this case are, in general, too complex to write down in closed form. However, some can be computed analytically, and perturbative unitarity bounds are imposed accordingly,
@@ V.5.1 Scattering Matrices
$M_{|0,0,0\rangle}=\left[\begin{array}[]{cccc}6\lambda_{11,11}&2\gamma_{11,11}&2\sqrt{2}\gamma_{11,12}&2\gamma_{11,22}\\ 2\gamma_{11,11}&12\beta_{11,11}&12\sqrt{2}\beta_{11,12}&12\beta_{11,22}\\ 2\sqrt{2}\gamma_{11,12}&12\sqrt{2}\beta_{11,12}&24\beta_{11,22}&12\sqrt{2}\beta_{12,22}\\ 2\gamma_{11,22}&12\beta_{11,22}&12\sqrt{2}\beta_{12,22}&12\beta_{22,22}\\ \end{array}\right].$ (81)
V.5.2 Perturbative Unitarity Bounds
Again, certain eigenvalues of the scattering matrices are too complicated to be expressed in closed form. However, others can be written analytically, and, for those, the corresponding perturbative unitarity bounds are then applied, as:
@@ V.5.2 Perturbative Unitarity Bounds
In order to compare our results with previous work, we consider the model studied in Ref. Mühlleitner et al. (2022), where the SM is extended by a neutral complex scalar singlet. In their notation, the quartic part of the scalar potential reads
@@
where $S=\frac{\chi_{1}+i\chi_{2}}{\sqrt{2}}$. Notice that the case with two real singlets in Eq. (78) is more general than the case with one complex singlet in Eq. (85). But, since the complex scalar singlet can be decomposed into two real scalar singlet fields, $\chi_{1}$ and $\chi_{2}$, we find that our results reproduce those of Mühlleitner et al. (2022), when we set $\gamma_{11,12}=\beta_{11,12}=\beta_{12,22}=0$, i.e. when we impose an additional $\mathbb{Z}_{2}$ that acts as $\chi_{2}\to-\chi_{2}$. [^5] Equivalently, one could choose $S=\frac{\chi_{2}+i\chi_{1}}{\sqrt{2}}$ and impose the $\mathbb{Z}_{2}$ symmetry on $\chi_{1}$. This procedure is generalizable to models with any number of complex neutral singlet scalars. By matching the terms in Eq. (85) to those in Eq. (78), we derive the mapping between the two notations that is provided in Table 6.
Table 6: Comparison of coupling notation.
Term Our Notation Notation in Mühlleitner et al. (2022)
@@ V.6.1 Scattering Matrices
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,12}&0\\ 2\lambda^{*}_{12,12}&2\lambda_{22,22}&0\\ 0&0&2(\lambda_{11,22}+\lambda_{12,21})\end{bmatrix}$, (89)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda_{12,12}\\ 0&0&2\lambda^{*}_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (90)
$M_{|0,0,0\rangle}=\left[\begin{array}[]{ccccc}6\lambda_{11,11}&2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&2\gamma_{11,11}&0&0\\ 2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&6\lambda_{22,22}&2\gamma_{22,11}&0&0\\ 2\gamma_{11,11}&2\gamma_{22,11}&12\beta_{11,11}&0&0\\ 0&0&0&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,12}\\ 0&0&0&6\lambda_{12,12}^{*}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)\\ \end{array}\right],$ (91)
@@ V.6.2 Perturbative Unitarity Bounds
$\left|\lambda_{11,11}+\lambda_{22,22}\pm\sqrt{(\lambda_{11,11}-\lambda_{22,22})^{2}+4|\lambda_{12,12}|^{2}}\right|$ $\leq 8\pi,$ (98)
$2\left|\lambda_{11,22}+2\lambda_{12,21}\pm 3|\lambda_{12,12}|^{2}\right|$ $\leq 8\pi.$ (99)
@@ V.7.1 Scattering Matrices
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\left[\begin{array}[]{cccc}2\gamma_{22,11}&2\gamma_{22,12}&0&0\\ 2\gamma_{22,12}&2\gamma_{22,22}&0&0\\ 0&0&2\gamma_{11,22}&2\gamma_{11,12}\\ 0&0&2\gamma_{11,12}&2\gamma_{11,11}\\ \end{array}\right],$ (103)
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,12}&0\\ 2\lambda_{12,12}^{*}&2\lambda_{22,22}&0\\ 0&0&2\lambda_{11,22}+2\lambda_{12,21}\end{bmatrix}$, (104)
$M_{|1,0,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\lambda_{12,21}&0&0\\ 2\lambda_{12,21}&2\lambda_{22,22}&0&0\\ 0&0&2\lambda_{11,22}&2\lambda_{12,12}\\ 0&0&2\lambda^{*}_{12,12}&2\lambda_{11,22}\end{bmatrix}$, (105)
@@
$A=\begin{bmatrix}24\,\beta_{11,22}&2\sqrt{2}\,\gamma_{11,12}&2\sqrt{2}\,\gamma_{22,12}&12\sqrt{2}\,\beta_{11,12}&12\sqrt{2}\,\beta_{12,22}\\ 2\sqrt{2}\,\gamma_{11,12}&6\,\lambda_{11,11}&2(2\,\lambda_{11,22}+\,\lambda_{12,21})&2\,\gamma_{11,11}&2\,\gamma_{11,22}\\ 2\sqrt{2}\,\gamma_{22,12}&2(2\,\lambda_{11,22}+\,\lambda_{12,21})&6\,\lambda_{22,22}&2\,\gamma_{22,11}&2\,\gamma_{22,22}\\ 12\sqrt{2}\,\beta_{11,12}&2\,\gamma_{11,11}&2\,\gamma_{22,11}&12\,\beta_{11,11}&12\,\beta_{11,22}\\ 12\sqrt{2}\,\beta_{12,22}&2\,\gamma_{11,22}&2\,\gamma_{22,22}&12\,\beta_{11,22}&12\,\beta_{22,22}\end{bmatrix}$, (108)
$B=\begin{bmatrix}2(\,\lambda_{11,22}+2\,\lambda_{12,21})&6\,\lambda_{12,12}\\ 6\,\lambda_{12,12}^{*}&2(\,\lambda_{11,22}+2\,\lambda_{12,21})\end{bmatrix}$. (109)
V.7.2 Perturbative Unitarity Bounds
@@
$\left|\gamma_{11,11}+\gamma_{11,22}\pm\sqrt{4\gamma_{11,12}^{2}+\left(\gamma_{11,11}-\gamma_{11,22}\right){}^{2}}\right|$ $\leq$ $8\pi,$ (115)
$\left|\gamma_{22,11}+\gamma_{22,22}\pm\sqrt{4\gamma_{22,12}^{2}+\left(\gamma_{22,11}-\gamma_{22,22}\right){}^{2}}\right|$ $\leq$ $8\pi.$ (116)
Table 7: Comparison of coupling notation.
Term Our Notation Notation of Boto et al. (2025)
@@
We verified our results against those reported in Boto et al. (2025) and found them to be in agreement. Table 7 summarizes the correspondence between the couplings used in our work and those of Boto et al. (2025).
@@ V.8.1 Scattering Matrices
$M_{|2,\frac{3}{2},\frac{1}{2}\rangle}=\begin{bmatrix}\delta_{11,11}&\delta_{12,11}\\ \delta_{12,11}^{*}&\delta_{22,11}\end{bmatrix}$, (119)
$M_{|1,\frac{1}{2},\frac{1}{2}\rangle}=\left[\begin{array}[]{cccc}2\gamma_{11,11}&2\gamma_{12,11}&0&2i\kappa_{12,11}^{*}\\ 2\gamma_{12,11}^{*}&2\gamma_{22,11}&-2i\kappa_{12,11}^{*}&0\\ 0&2i\kappa_{12,11}&\delta_{11,11}&\delta_{12,11}^{*}\\ -2i\kappa_{12,11}&0&\delta_{12,11}&\delta_{22,11}\\ \end{array}\right],$ (120)
$M_{|1,1,0\rangle}=\begin{bmatrix}2(\lambda_{11,22}-\lambda_{12,21})&2i\sqrt{2}\kappa^{*}_{12,11}\\ -2i\sqrt{2}\kappa_{12,11}&2\zeta_{11,11}\end{bmatrix}$, (121)
$M_{|2,1,1\rangle}=\begin{bmatrix}2\lambda_{11,11}&2\sqrt{2}\lambda_{11,12}&2\lambda_{12,12}\\ 2\sqrt{2}\lambda_{11,12}^{*}&2(\lambda_{11,22}+\lambda_{12,21})&2\sqrt{2}\lambda_{12,22}\\ 2\lambda_{12,12}^{*}&2\sqrt{2}\lambda_{12,22}^{*}&2\lambda_{22,22}\end{bmatrix}$, (122)
$M_{|1,0,1\rangle}=\left[\begin{array}[]{cccc}2\lambda_{11,11}&2\lambda_{11,12}^{*}&2\lambda_{11,12}&2\lambda_{12,21}\\ 2\lambda_{11,12}&2\lambda_{11,22}&2\lambda_{12,12}&2\lambda_{12,22}\\ 2\lambda_{11,12}^{*}&2\lambda_{12,12}^{*}&2\lambda_{11,22}&2\lambda_{12,22}^{*}\\ 2\lambda_{12,21}&2\lambda_{12,22}^{*}&2\lambda_{12,22}&2\lambda_{22,22}\\ \end{array}\right],$ (123)
$M_{|0,0,0\rangle}=\left[\begin{array}[]{cccccc}6\lambda_{11,11}&6\lambda_{11,12}^{*}&6\lambda_{11,12}&2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&\sqrt{2}\delta_{11,11}&2\gamma_{11,11}\\ 6\lambda_{11,12}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,12}&6\lambda_{12,22}&\sqrt{2}\delta_{12,11}&2\gamma_{12,11}\\ 6\lambda_{11,12}^{*}&6\lambda_{12,12}^{*}&2\left(\lambda_{11,22}+2\lambda_{12,21}\right)&6\lambda_{12,22}^{*}&\sqrt{2}\delta_{12,11}^{*}&2\gamma_{12,11}^{*}\\ 2\left(2\lambda_{11,22}+\lambda_{12,21}\right)&6\lambda_{12,22}^{*}&6\lambda_{12,22}&6\lambda_{22,22}&\sqrt{2}\delta_{22,11}&2\gamma_{22,11}\\ \sqrt{2}\delta_{11,11}&\sqrt{2}\delta_{12,11}^{*}&\sqrt{2}\delta_{12,11}&\sqrt{2}\delta_{22,11}&4\alpha_{11,11}&\sqrt{2}\zeta_{11,11}\\ 2\gamma_{11,11}&2\gamma_{12,11}^{*}&2\gamma_{12,11}&2\gamma_{22,11}&\sqrt{2}\zeta_{11,11}&12\beta_{11,11}\\ \end{array}\right].$ (124)
V.8.2 Perturbative Unitarity Bounds
Since certain eigenvalues are too complex to evaluate analytically, the perturbative unitarity constraints are shown only for the remaining eigenvalues of the zero partial-wave amplitude matrix. These read
$\left|\zeta_{11,11}+\lambda_{11,22}-\lambda_{12,21}\pm\sqrt{\left(-\zeta_{11,11}-\lambda_{11,22}+\lambda_{12,21}\right)^{2}+8|\kappa_{12,11}|^{2}}\right|$ $\leq 8\pi,$ (125)